40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8921     Entry time: Thu Jul 25 02:53:00 2013
Author: Koji 
Type: Update 
Category: General 
Subject: IFO Alignment after TT flipping - no progress 

There was no progress tonight after Jenne left.
I could not find any reasonable fringes of the IFO after 3 hours of optics jiggling.

* I jiggled TT1 and TT2. The slider has not been restored.
We should probably look at the value in the day time and revert them.
(Still this does not ensure the recovery of the previous pointing because of the hysteresis)

* The arms are still aligned for the green.
It's not TEM00 any more because of the vent/drift but the fringe is visible (i.e. eigenaxis is on the mirror)

* As we touched PR3, the input pointing is totally misaligned.

To Do / Plan

* We need to find the resonance of the yarm by the input TTs. Once the resonance is found, we will align the PRM.

* Move the BS to find the xarm resonance.

* Finally align SRM

* It was not possible to find the resonance of the yarm without going into the chamber. Definitely we can find the spot on the ITMY by a card, but we are not sure the beam can hit the ETMY. And the baffles makes the work difficult.

* One possibility is to align the input beam so that the ITMY beam is retroreflected to the PRM. I tried it but the beam was not visible form the camera.

ELOG V3.1.3-